# Evaluating Time-Dependent Methods and Seasonal Effects in Code Technical Debt Prediction

### v1

@@ Abstract
Background. Code Technical Debt prediction has become a popular research niche in recent software engineering literature. Technical Debt is an important metric in software projects as it measures the effort professionals need to clean the code. Therefore, predicting its future behaviour becomes a crucial task. However, there is no well-defined and consistent approach which is able to completely capture the features that impact on the evolution of Code Technical Debt.
Objective. The goal of this study is to evaluate the impact of considering time-dependent techniques as well as seasonal effects in temporal data in the prediction performance within the context of Code Technical Debt. The study adopts already existing, yet not extensively adopted, time-dependent prediction techniques and compares their prediction performance to that of commonly used Machine Learning models. Further, the study strengthens the evaluation of time-dependent methods by extending the analysis to capture the impact of seasonality in Code Technical Debt data.
Methods. We trained 11 prediction models for using the commit history of 31 open-source projects developed with Java. We predicted the future observations of the SQALE index to evaluate their predictive performance.
Results. Our study confirms the positive impact of considering time-dependent techniques. The adopted multivariate time series analysis model ARIMAX overcame the rest of the adopted models. Incorporating seasonal effects led to an enhancement in the predictive performance of the adopted time-dependent techniques. However, the impact of this effect was found to be relatively modest.
Conclusions. The findings of this study corroborate our position in favour of implementing techniques that capture the existing time dependence within historical data of software metrics, specifically in the context of this study, namely, Code Technical Debt. This necessitates the utilisation of techniques that can effectively address this evidence.
@@
In Software Engineering (SE), and especially in Mining Software Repository (MSR) studies, researchers often investigate the relationships among different variables collected from the history of software projects. As an example, researchers have investigated correlations between two variables such as code smells [1, 2, 3], their trend over time [4], and the impact of different software qualities [5, 6, 7]. However, a large number of these studies have omitted the potential impact hidden on the temporal dependency of the analyzed variables and the possible threats related to the usage of statistical techniques not designed for analyzing temporally dependent data [8]. As an example, the introduction of a technical issue in a commit heavily depends on the code that was present in the repository before the commit. Still, most studies have not considered this aspect, mainly due to a lack of clear guidelines on the statistical techniques that can be used in this context. Saarimäki et al. [8] highlighted three main issues in previous studies: 1) discarding the temporal nature of the commits, 2) assuming independence of data, and 3) mixing projects of different sizes, where big projects overwhelm small ones. To overcome these issues, they proposed to analyze dependent data in the SE field, considering the dependency through the time effect [8]. Within the provided study context, we leverage the analysis of the Code Technical Debt (TD) prediction to evaluate the impact of the time-dependence factor. Code TD is an important metric in software projects as it measures the effort professionals need to put into making the code clean. Therefore, Code TD denotes time dependence on past mistakes. Consequently, data analysis techniques that assume this dependence should be considered.
@@
In this direction, the research community traditionally used Machine Learning (ML) models [9, 10]. Similarly, we identified two relevant works that applied Time-Series Analysis (TSA) to predict TD [11, 12] comparing univariate and multivariate TSA models. The obtained results were promising in terms of prediction performance. Hence, we aim to analyze the robustness of TSA techniques predicting time-dependent variables such as Code TD by comparing them with traditional ML methods. Therefore, we aim to identify existing potential variables that can help explain the behavior of Code TD.
@@
We designed and conducted an empirical study to evaluate the Code TD prediction performance of different TSA models and a set of time-agnostic ML models commonly used in the literature, which do not take into consideration the temporal factor of the data. Furthermore, we assessed the impact of the seasonality factor in the observed data. For that, we adopted the seasonally adjusted version of the TSA models already used in SE literature and measured their performance to quantify the impact of addressing the seasonality factor. As the adopted potential variable on the description of the Code TD, we considered the SQALE index (also called TD) computed by SonarQube (SQ) to depict TD. Specifically, we adopted two different multivariate applications of the Autoregressive Integrated Moving Average (ARIMA) model [13] as well as their modified counterparts which control the seasonality component of the data, and seven ML models to perform the confronted evaluation (four linear and three non-linear). Additionally, we implemented a backward variable selection approach [14, 15] to provide the models with the best variable combination to help improve their efficiency.
@@
Our results revealed a robust superiority obtained by the multivariate TSA models and demonstrated a slight improvement in the predictive performance when adjusting the models for capturing the seasonality effect. Especially, the Seasonally adjusted Autoregressive Integrated Moving Average with extra regressors (SARIMAX) and its unadjusted counterpart ARIMAX resulted in being the best prediction models among the methods considered in this study, both the mentioned methods being the multivariate application of the commonly known ARIMA model. However, the ARIMAX model demonstrated to be the model robustly overtaking the rest of ML algorithms in terms of predictive performance.
Therefore, the main contributions of this paper are:
@@
An improvement to the current knowledge of Code TD prediction with an emphasis on investigating the importance of including time-dependence factors in the prediction.
@@
A large-scale comparison among the time-dependent prediction models and ML prediction models already adopted in the literature.
@@
Section 2 describes the background our paper is based on. In Section 3, we present the empirical study design. In Section 4, we present the obtained results, and Section 5 shows the derived discussion based on them. Section 6 describes the identified threats to the validity of our study. Section 7 presents related work, and Section 8 divulges the conclusions of this study and outlines the potential future work.
@@ 2 Background
SQ is one of the most common static analysis tools adopted both in academia [16, 17] and in industry [18].
It aims to help developers create more clean, secure, maintainable, readable, and modular code. SQ evaluates the code by testing it against a predefined set of rules. The rules are considered coding standards and breaking them is deemed undesirable. Therefore, each time a piece of code violates a rule, SQ generates an issue presented to the developers.
SQ can be accessed as a service through the sonarcloud.io platform or downloaded and run on a private server. The tool supports 29 different programming languages including Java, Python, C++, and JavaScript.111https://docs.SQ.org/9.6/analyzing-source-code/languages/overview/ It performs various calculations, including measuring metrics like lines of code and code complexity. It also establishes specific thresholds known as “quality gates” for each metric and rule.
In this study, the selected dataset (Section 3.2) analyzed the projects with SQ version 7.5.
This version includes three rule categories:
@@
Maintainability rules named as code smells that decrease code readability and code modifiability. It is important to note that the term “code smells” adopted in SQ contains only a subset of the well-known code smells defined by Fowler et al. [19].
Security rules named as Vulnerability that creates a problem that impacts the application’s security.
SQ internally uses the SQALE methodology [20] to compute the technical debt ratio (TD ratio) and to classify the project to an SQALE rating considering the time to fix all open issues classified as code smell. Moreover, SQ computes other two types of debt called reliability remediation effort and security remediation effort, based on respectively on the time to fix all the open issues classified as bugs and as vulnerabilities.
@@ 3 The Empirical Study Design
We now describe our study reporting the goal and research questions, the context, data collection, and data analysis. We designed our experimental study based on the guidelines defined by Wohlin et al. [21].
@@ 3.1 Goal and Research Questions
Based on this goal, our formulated Research Questions (RQs) are presented as follows:
@@
The defined RQ1 compares two different multivariate TSA approaches already implemented in the SE literature to determine which approach can predict better Code TD. Evaluating the prediction performance of the latest models used to date that control the time-dependence factor within the data enables us to determine which of these models should be considered by practitioners when deliberating which prediction techniques to use when predicting Code TD. To address RQ1, we collected as Code TD the SQALE index metric computed by SQ (as explained in Section 2), as well as Code Smell rule violations generated by SQ as potential explainers of the SQALE index’s behavior.
The first approach, referred to as ARIMA + LM, combines the use of multiple ARIMA models and linear regression models (LM) [12]. This approach utilizes a univariate ARIMA model for the prediction of each independent variable and utilizes the predicted values inside a regression model to predict future values of the dependent variable. The second approach, referred to as ARIMAX consists of constructing a multivariate ARIMA model for predicting values of the dependent variable [11]. In the ARIMAX model, the history of the dependent variable as well as the past values of the considered independent variables are used as features to train the model and provide future predictions of the dependent variable.
@@
The goal of this study is not only to establish the applicability of the most efficient time-dependent techniques for Code TD prediction but also to evaluate how the adopted techniques perform against other prediction techniques commonly used in the SE literature. These techniques are mostly ML-based algorithms that do not consider the temporal factor when providing future predictions. Therefore, we formulate our RQ2:
@@
To assess the applicability of a time-dependent approach for Code TD prediction (RQ2), in our case multivariate TSA, we need to compare its performance with other approaches that do not take into consideration the time dependency of the data. In this study, the comparison is made with several ML models.
Similarly, as for RQ1, our prior assumption is that the prediction performance of multivariate TSA methods for Code TD prediction overcomes that of ML models. Moreover, TSA methods are based on temporally ordered serialized data, which helps the models capture potential temporal factors, for instance, the seasonality of the observation [23]. Models can be adjusted to these factors to observe their impact on the predicted results. Therefore, only comparing already used TSA methods is not sufficient and we formulate a third research question as follows:
@@
As explained, RQ3 expands RQ1. For that, given the modification in the models to handle seasonality patterns, RQ3 compares the new models with the former ones in terms of Code TD prediction and therefore evaluates the impact of seasonality in the prediction performance.
Our prior assumption is that SARIMAX, SARIMA + LM approaches and the seasonally adjusted models of the previously defined TSA model counterparts perform better than their former model format which did not handle seasonality patterns. Further, we leverage the comparison proposed in RQ1, thus we assume that using the multivariate SARIMAX model is more informative than the SARIMA+LM model, given the extension of both TSA models to capture the impact of the seasonality effects.
Similarly, evaluating the seasonality adjustment in the prediction performance compared to the selected ML algorithms becomes paramount to providing a complete contribution to the SE community through this study. Therefore, we ask:
@@
For all RQs, the models were evaluated using the following set of performance metrics: Mean Absolute Percentage Error (MAPE), Mean Absolute Error (MAE), and Root Mean Square Error (RMSE) which will be further described in Section 3.5.5.
@@ 3.2 Context
As the context of our study, we consider the Technical Debt Dataset [24] version 2.0 adopted by Saarimäki et al. [8]. The data set contains 31 Java projects from the Apache Software Foundation (ASF) repository.222http://apache.org The projects in the data set were selected based on “criterion sampling” [25], that fulfil all of the following criteria: developed in Java, older than three years, more than 500 commits and 100 classes, and usage of an issue tracking system with at least 100 issues reported. The projects were selected, maximizing their diversity and representation by considering a comparable number of projects concerning project age, size, and domain. The projects can be considered mature, due to the strict review and inclusion process required by the ASF. Moreover, the included projects regularly review their code and follow a strict quality process.333https://incubator.apache.org/policy/process.html Full details of the data can be found in the online repository of the data set.444https://github.com/clowee/The-Technical-Debt-Dataset We adopted this dataset since all its projects were already analyzed with SQ and it contains the data related to the SQALE Index.
@@ 3.3 Variables
In this section, we describe the dependent and independent variables used in the adopted models for Code TD prediction:
Dependent variable: We considered the metric SQALE index computed by SQ. The SQALE index denotes the efforts to fix all Code Smell rules violated in the code and is measured in minutes. SQ denotes TD using the total value of the SQALE index. The metric is continuous and, therefore, is suitable for the TSA models adopted in this study. Further, it provides a clear definition of Code TD to answer the goal of our study.
Independent variables: We considered the Code Smell rule violations detected by SQ. Code smell rule violations describe the number of violations for each SQ rule of code smell type. The TD dataset has issues from 205 different rules of type code smell and each rule is included as a separate independent variable in the study. The nature of the collected issues is discrete, as it is founded on the count of the issues detected in each SQ analysis for each of the Code Smell types.

### v2

@@ Abstract
Background. Code Technical Debt (Code TD) prediction has gained significant attention in recent software engineering research. However, no standardized approach to Code TD prediction fully captures the factors influencing its evolution.
Objective. Our study aims to assess the impact of time-dependent models and seasonal effects on Code TD prediction. It evaluates such models against widely used Machine Learning models also considering the influence of seasonality on prediction performance.
Methods. We trained 11 prediction models with 31 Java open-source projects. To assess their performance, we predicted future observations of the SQALE index. To evaluate the practical usability of our TD forecasting model and their impact on practitioners, we surveyed 23 software engineering professionals.
Results. Our study confirms the benefits of time-dependent techniques, with the ARIMAX model outperforming the others. Seasonal effects improved predictive performance, though the impact remained modest. ARIMAX/SARIMAX models demonstrated to provide well-balanced long-term forecasts. The survey highlighted strong industry interest in short- to medium-term TD forecasts.
Conclusions. Our findings support using techniques that capture time dependence in historical software metric data, particularly for Code TD. Effectively addressing this evidence requires adopting methods that account for temporal patterns.
@@
In Software Engineering (SE), and specially in Mining Software Repository (MSR) studies, researchers often investigate the relationships among different variables collected from the history of software projects. As an example, researchers have investigated the correlations between two variables, such as code smells [1, 2, 3], their trend over time [4], and the impact of different qualities of software [5, 6, 7]. However, many of these studies have omitted the hidden potential impact on the temporal dependency of the variables analyzed and the possible threats related to statistical techniques not designed for analyzing temporally dependent data [8]. As an example, the introduction of a technical issue in a commit heavily depends on the code that was present in the repository before said commit. However, most studies have not considered this aspect, mainly due to a lack of clear guidelines on the statistical techniques used in this context. Saarimäki et al. [8] highlighted three main issues in previous studies: 1) discarding the temporal nature of the commits, 2) assuming independence of data, and 3) mixing projects of different sizes, where big projects overwhelm small ones. To overcome these issues, they proposed to analyze the dependent data in the SE field, considering the dependency through the time effect [8]. Within the provided study context, we use the code technical debt (code TD) prediction analysis to evaluate the impact of the time dependence factor. Code TD is an essential metric in software projects as it measures professionals’ efforts to clean the code. Therefore, code TD denotes the time dependence on past mistakes. Consequently, data analysis techniques that assume this dependence should be considered.
@@
In this direction, the research community traditionally uses machine learning (ML) models [9, 10]. Similarly, we identified two relevant works that applied Time Series Analysis (TSA) to predict code TD [11, 12] comparing univariate and multivariate TSA models. The results obtained were promising in terms of prediction performance. Hence, our objective is to analyze the robustness of TSA techniques predicting time-dependent variables such as code TD by comparing them with traditional ML methods. Therefore, we are striving to identify existing potential variables that can help explain the behavior of code TD.
@@
We designed and conducted an empirical study to evaluate the code TD prediction performance of different TSA models and a set of time-agnostic ML models commonly used in the literature, which do not consider the data’s temporal factor. Furthermore, we assessed the impact of the seasonality factor on the observed data. For that, we adopted the seasonally adjusted version of the TSA models already used in the SE literature and measured their performance to quantify the impact of addressing the seasonality factor. As the potential dependent variable adopted for the description of the code TD, we considered the SQALE index metric, computed by SonarQube (SQ) to measure code TD. Specifically, we adopted two different multivariate applications of the Autoregressive Integrated Moving Average (ARIMA) model [13] as well as their modified counterparts, which control the seasonality component of the data, and seven ML models to perform the confronted evaluation (four linear and three non-linear). Furthermore, we implemented a backward variable selection approach [14, 15] to provide models with the best combination of variables to help improve their efficiency. The best resulting multivariate TSA models were afterwards used to explore their accuracy when performing long-term time-series forecasting, a facet of interest in software maintenance activities [10].
Finally, to evaluate the practical usability of our TD forecasting model and their impact on practitioners, we surveyed 23 software engineering professionals. The survey was used to determine their willingness to use the model, the usability of the model in real-world TD management, and the predicted window of choice for decision making. The survey asked questions about the professional experience of the respondents, their knowledge of TD and whether they manage TD issues in real life. In addition, the participants provided information on the best prediction horizons, trading off precision for predictions with practicability. The study supports the notion that practitioners desire high-precision predictions in short to medium-term horizons, which is consistent with agile software development cycles and allows for the use of TD predictions within industry pipelines.
@@
Our results revealed a robust superiority obtained by the multivariate TSA models and demonstrated a slight improvement in predictive performance when adjusting the models to capture the seasonality effect. Basically, Seasonally adjusted Autoregressive Integrated Moving Average with extra regressors (SARIMAX) [16] and its non-adjusted counterpart ARIMAX [17] were the best prediction models among the methods considered in this study. These both methods implement a multivariate variant of the commonly known ARIMA model. However, the ARIMAX model outperforms the other ML algorithms in predictive performance [17]. Long-term forecast performance showed increased MAPE errors over time but remained balanced across projects. The MAPE error of the SARIMAX model increased from 1.08% at 6 months to 4.19% at 36 months, while ARIMAX remained at 3.15% after 12 biweekly periods and 5.97% after 72 periods. The variance remained low, with stable trends in descriptive statistics. Finally, our survey reveals good industry confidence in our tool and its effectiveness for short- to medium-term forecasting windows, which aligns closely with industry needs.
The main contributions of this paper are:
@@
An improvement in current knowledge of the prediction of the code TD with an emphasis on investigating the importance of including time-dependent factors.
@@
A large-scale comparison among the time-dependent and ML prediction models already adopted in the literature.
A demonstration of the promising capabilities of classic, yet efficient multivariate TSA forecasting models on long-term code TD forecasting, aiming to enable system engineers and project managers to perform long-term effective software maintenance.
An industrial survey with 23 experts confirming strong industry interest in short- to medium-term TD forecasts, aligning with agile workflows.
@@
Section 2 describes the background on which our paper is based. In Section 3, we present the empirical study design. Section 4 presents the results obtained, and Section 5 shows the derived discussion based on them. Section 6 describes the identified threats to the validity of our study. Section 7 presents related work and Section 8 divulges the conclusions of this study and outlines the potential future work.
@@ 2 Background
SQ is one of the most commonly adopted static analysis tools both in academia [18, 19] and in industry [20]. We aim to help developers create more clean, secure, maintainable, readable, and modular code. SQ evaluates the code by testing it against a predefined set of rules. The rules are considered coding standards, and breaking them is deemed undesirable. Therefore, each time a code violates a rule, the SQ presents an issue to the developers. Practitioners can use SQ on its official ”as a Service” flavor hosted on the sonarcloud.io website or the on-demand version to run on a private server. The tool supports 29 programming languages: Java, Python, C++, and JavaScript.111https://docs.SQ.org/9.6/analyzing-source-code/languages/overview/ It performs various calculations, including measuring metrics like lines of code and code complexity. It also establishes specific thresholds known as “quality gates” for each metric and rule.
In this study, the selected dataset (Section 3.2) analyzed projects with SQ version 7.5, which includes three rule categories:
@@
Maintainability rules named as code smells that decrease code readability and modifiability. It is important to note that the term “code smells” adopted in SQ contains only a subset of the well-known code smells defined by Fowler et al. [21].
Security rules named as Vulnerability, creates a problem that impacts the application’s security.
To quantify and therefore measure the existing technical debt in software projects, SQ leverages a series of metrics, mostly quantitative translations of code analysis metrics. In the version considered for this study, SQ computed three types of TD, reliability remediation effort and security remediation effort, which consisted on the time to fix all the open issues classified as bugs and vulnerabilities, and the technical debt ratio, which SQ named as SQALE index resembling the naming of SQALE methodology [22] used in older versions, and consisting on the time to fix all open issues classified as code smell.
@@ 3 The Empirical Study Design
We now describe our study, reporting the goal and research questions, the context, the collection and analysis of data. We designed our experimental study based on the guidelines defined by Wohlin et al. [23].
@@ 3.1 Goal and Research Questions
Based on this goal, our formulated Research Questions (RQs) are presented next.
@@
RQ1 compares two different multivariate TSA approaches already implemented in the SE literature to determine which approach can predict code TD better. Evaluating the prediction performance of the latest models that control the time dependence factor within the data enables us to determine which of these models should be considered by practitioners when deliberating which prediction techniques to use when predicting code TD. To address RQ1, we collected as code TD the SQALE index metric computed by SQ (as explained in Section 2), and Code Smell rule violations generated by SQ as potential explanations of the SQALE index’s behavior.
The first approach, known as ARIMA + LM, combines multiple ARIMA models and linear regression models (LM) [12]. This approach utilizes a univariate ARIMA model for the prediction of each independent variable. Using predicted values within a regression model, it predicts future values of the dependent variable. The second approach, known as ARIMAX, consists of building a multivariate ARIMA model to predict the values of the dependent variable [11]. In the ARIMAX model, the history of the dependent variable and the past values of the considered independent variables are used as features to train the model and provide future predictions of the dependent variable.
@@
The purpose of this study is to establish the applicability of the most efficient time-dependent techniques for code TD prediction and evaluate how the adopted techniques perform against other prediction techniques commonly used in the SE literature. These techniques are mostly ML-based algorithms that do not consider the temporal factor when providing future predictions. Therefore, we formulate our RQ2.
@@
To assess the applicability of a time-dependent approach for code TD prediction (RQ2), in our case multivariate TSA, we need to compare its performance with other approaches that do not consider the data’s time dependency. In this study, the comparison is made with several ML models.
Similarly, as for RQ1, our prior assumption is that the prediction performance of multivariate TSA methods for code TD prediction overcomes that of ML models. Moreover, TSA methods are based on temporally ordered serialized data, helping the models capture potential temporal factors, for example, the seasonality of the observation [25]. Models can be adjusted to these factors to observe their impact on the predicted results. Therefore, only comparing already used TSA methods is not sufficient, and we formulate a third research question.
@@
As explained, RQ3 expands RQ1. For that, given the modification of the models to handle seasonality patterns, RQ3 compares the new models with the former in terms of the prediction of the code TD and therefore evaluates the impact of seasonality in the prediction performance.
Our prior assumption is that SARIMAX, SARIMA + LM approaches and seasonally adjusted models of the previously defined TSA model counterparts perform better than their former model format, which did not handle seasonality patterns. In addition, we take advantage of the comparison proposed in RQ1. Thus we assume that using the multivariate SARIMAX model is more informative than the SARIMA+LM model, given the extension of both TSA models to capture the impact of the seasonality effects.
Similarly, evaluating the seasonality adjustment in prediction performance compared to the selected ML algorithms becomes paramount to contributing to the SE community through this study. Therefore, we ask a fourth question.
@@
For all RQs described so far, we evaluated the models using three key performance metrics: Mean Absolute Percentage Error (MAPE), Mean Absolute Error (MAE), and Root Mean Square Error (RMSE) (see Section 3.6.5 for a detailed description of these metrics).
RQ5. How accurate are multivariate time-dependent methods on long-term forecasting?
A challenge that emerges from the concept of code TD prediction is the need to make multi-step forecasts, that is, forecasts for more than one time-step into the future. Providing reasonable long-term forecasts would allow system engineers and project managers to perform long-term effective software maintenance. While single-step predictions can be useful for short-term planning, multi-step forecasting offers more practical insights for software maintenance and allows project managers to make strategic decisions for the future. Therefore, through the defined RQ5 we will investigate the performance of the resulting best TSA models when performing multi-step forecasting to enable practitioners observe the long-term impact of their current project maintenance level within the context of code TD.
RQ6. How valuable do software practitioners perceive a Technical Debt forecasting to be? • RQ6.1. What forecasting window combinations do practitioners consider acceptable and useful for Technical Debt prediction? • RQ6.2. What is the preferred scope of Technical Debt prediction among practitioners?
While many techniques exist to measure and manage code TD retrospectively, forecasting its evolution could allow practitioners to take proactive measures. Understanding how practitioners perceive code TD forecasting is crucial to determining whether such predictive models can influence decision-making processes in software projects or are just academic interests (RQ6). If practitioners do not see substantial value, efforts in this direction might need re-evaluation or refinement to align with industry needs.
To further explore this aspect, we investigate the specific forecasting configurations that practitioners deem useful. To effectively forecast the evolution of code TD (RQ6.1), it is crucial to determine the appropriate time horizon for predictions. Practitioners may prefer short-term forecasts over weeks or months to aid in immediate refactoring efforts. In contrast, long-term forecasts, spanning years, could be more useful for strategic planning. Knowing which time frames practitioners consider valid and useful ensures that code TD prediction models meet actual industry needs, thus facilitating their practical application. Moreover, selecting suitable forecast windows is just one part of the equation. Defining the preferred scope of code TD predictions to customize these forecasts properly is equally important.
Code TD can appear at different levels, from specific code components to software projects. Some practitioners focus on predicting TD at the module or class level to aid in detailed refactoring efforts (RQ6.2). In contrast, others prefer predictions at the system-wide level to help with architectural decisions. Understanding the preferred scope ensures that code TD forecasting tools provide valuable insights at the appropriate level of abstraction, making them more useful in practitioners’ daily workflows. Our survey offers a comprehensive view on creating Technical Debt forecasting methods that are practical and actionable for software professionals.
@@ 3.2 Context
In the context of our study, we consider the Technical Debt Dataset [28] version 2.0 adopted by Saarimäki et al. [8]. The data set contains 31 Java projects from the Apache Software Foundation (ASF) repository.222http://apache.org The projects in the data set were selected based on “criterion sampling” [29], requiring all of the following criteria: developed in Java, older than 3 years, more than 500 commits and 100 classes, and the usage of an issue tracking system with at least 100 issues reported. The projects were selected, maximizing their diversity and representation by considering a comparable number of projects on the age, size, and domain of the project. The projects can be considered mature due to the strict review and inclusion process required by the ASF. In addition, the included projects regularly review their code and follow a strict quality process.333https://incubator.apache.org/policy/process.html Full details of the data can be found in the online repository of the data set.444https://github.com/clowee/The-Technical-Debt-Dataset We adopted this data set since all its projects were already analyzed with SQ and it contains the data related to the SQALE index.
@@ 3.3 Variables
In this section, we describe the dependent and independent variables used in the models adopted for the prediction of code TD.
Dependent variable: We consider the metric SQALE index computed by SQ. The SQALE index denotes the efforts to fix all the Code Smell rules that are violated in the code and is measured in minutes. SQ denotes TD using the total value of the SQALE index. The metric is continuous and, therefore, is suitable for the TSA models adopted in this study. In addition, it provides a clear definition of code TD to answer the objective of our study.
Independent variables: We consider the Code Smell rule violations detected by SQ. Code smell rule violations describe the number of violations for each SQ rule of code smell type. The TD dataset has issues from 205 different rules of type code smell, and each rule is included as a separate independent variable in the study. The nature of the collected issues is discrete as it is based on the count of issues detected in each SQ analysis for each code smell type.